How many cubic centimeters (cc)of a 25% solution should be added to 10 cubic centimeters of a 60% solution in order to get a 30% solution?
Answer by haileytucki(390)

x(0.25)+10(0.6)=(x+10)(0.3)
Since 10(0.6) does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 10(0.6) from both sides.
x(0.25)=-10(0.6)+(x+10)(0.3)
Simplify the right-hand side of the equation.
x(0.25)=0.3x-3
Multiply x by each term inside the parentheses.
0.25x=0.3x-3
Since 0.3x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 0.3x from both sides.
0.25x-0.3x=-3
Since 0.25x and -0.3x are like terms, add -0.3x to 0.25x to get -0.05x.
-0.05x=-3
Divide each term in the equation by -0.05.
-(0.05x)/(-0.05)=-(3)/(-0.05)
Cancel the common factor of -0.05.
x=-(3)/(-0.05)
Simplify the right-hand side of the equation by simplifying each term.
x=60 (cubic cenimeters)